let V be a vector space of 3x3 symmetric matric.show that if A,B in V, then <A,B>=tr(AB) in inner product space


1
Expert's answer
2021-01-13T18:20:36-0500

Let


"A=\begin{bmatrix}
a_{11} & a_{12} & a_{13}\\
a_{21} & a_{22} & a_{23}\\
a_{31} &a_{32} & a_{33}
\end{bmatrix}" and "B=\begin{bmatrix}
b_{11} & b_{12} & b_{13}\\
b_{21} & b_{22} & b_{23}\\
b_{31} &b_{32} & b_{33}
\end{bmatrix}"


"\Rightarrow AB=\begin{bmatrix}
a_{11} & a_{12} & a_{13}\\
a_{21} & a_{22} & a_{23}\\
a_{31} &a_{32} & a_{33}
\end{bmatrix}\begin{bmatrix}
b_{11} & b_{12} & b_{13}\\
b_{21} & b_{22} & b_{23}\\
b_{31} &b_{32} & b_{33}
\end{bmatrix}"


"(AB) =\begin{bmatrix}
a_{11}b_{11} +a_{21}b_{12} +a_{31}b_{13} & a_{11}b_{12} + a_{12}b_{22}+a_{13}b_{32}& a_{11}b_{13} + a_{12}b_{23}+a_{13}b_{33}\\
 a_{21}b_{11} + a_{22}b_{21}+a_{23}b_{31}& a_{21}b_{12} + a_{22}b_{22}+a_{23}b_{32} & a_{21}b_{13} + a_{22}b_{23}+a_{23}b_{33}\\
a_{31} b_{11} + a_{32}b_{21}+ a_{33}b_{31}&a_{31} b_{12} + a_{32}b_{22}+ a_{33}b_{32}&a_{31} b_{13} + a_{32}b_{23}+ a_{33}b_{33}
\end{bmatrix}"


"(AB)^T =\begin{bmatrix}
a_{11}b_{11} +a_{21}b_{12} +a_{31}b_{13} & a_{21}b_{11} + a_{22}b_{21}+a_{23}b_{31}& a_{31} b_{11} + a_{32}b_{21}+ a_{33}b_{31}\\
 a_{11}b_{12} + a_{12}b_{22}+a_{13}b_{32}& a_{21}b_{12} + a_{22}b_{22}+a_{23}b_{32} &a_{31} b_{12} + a_{32}b_{22}+ a_{33}b_{32} \\
a_{11}b_{13} + a_{12}b_{23}+a_{13}b_{33}&a_{21}b_{13} + a_{22}b_{23}+a_{23}b_{33}&a_{31} b_{13} + a_{32}b_{23}+ a_{33}b_{33}
\end{bmatrix}"


"<A ,B>=a_{11}b_{11} +a_{21}b_{12} +a_{31}b_{13} + a_{21}b_{11} + a_{22}b_{21}+a_{23}b_{31}+ a_{31} b_{11} + a_{32}b_{21}+ a_{33}b_{31}\n a_{11}b_{12} + a_{12}b_{22}+a_{13}b_{32}+ a_{21}b_{12} + a_{22}b_{22}+a_{23}b_{32} +a_{31} b_{12} + a_{32}b_{22}+ a_{33}b_{32}\na_{11}b_{13} + a_{12}b_{23}+a_{13}b_{33}+a_{21}b_{13} + a_{22}b_{23}+a_{23}b_{33}+a_{31} b_{13} + a_{32}b_{23}+ a_{33}b_{33}"


"=\Sigma_{i=1}^{i=3}\Sigma_{j=1}^{j=3}a_{ij}b_{ij}"